OpenStudy (anonymous):

Estimate each quantity in terms of powers of ten. 2^70 4^90

OpenStudy (anonymous):

we know that 2^10 = 1024 which is approximately equal to 10^3 so 2^70 = (2^10)7 = (10^3)^7 = 10^21

OpenStudy (anonymous):

these are of course approximate values

OpenStudy (anonymous):

The question is to estimate, after all.

OpenStudy (anonymous):

4^90 = (4^5)^18 = (2^10)^18 2^10, as dhatraditya said, is 1024, which is roughly 10^3 = (10^3)^18 = 10^54